# Physics-Conditioned Diffusion Models for Lattice Gauge Theory

### v1

@@ Appendix A Langevin Sampler
4:     $\alpha_{i}\leftarrow\epsilon\cdot g\cdot t_{i}^{2}/t_{N_{T}}^{2}$triangleright $\alpha_{i}$ is the step size.
@@
4:     $\alpha_{i}\leftarrow\epsilon\cdot g\cdot t_{i}^{2}/t_{N_{T}}^{2}$triangleright $\alpha_{i}$ is the step size.
@@
9:         if $t_{i}<0.01$ then
10:              $P_{i}\leftarrow\min\left\{1,\frac{p(\tilde{\psi}_{\tau+1})q(\tilde{\phi}_{\tau}|\tilde{\psi}_{\tau+1})}{p(\tilde{\phi}_{\tau})q(\tilde{\psi}_{\tau+1}|{\tilde{\phi}_{\tau}})}\right\}$
@@ Appendix B Autocorrelation Time
Although in our approach multiple Markov chains are employed to generate independent samples, the dynamics in a single chain can be compared with other Monte Carlo methods. For traditional Monte Carlo methods, the characteristic length of autocorrelations for any observable $\mathcal{O}$ can be defined by the integrated autocorrelation time $\tau^{\mathrm{int}}_{\mathcal{O}}$ Madras:1988ei .
@@
Fig. 7 compares the integrated autocorrelation time of the topological charge, $\tau^{\mathrm{int}}_{Q}$, for HMC and DM. For both methods, $\tau^{\mathrm{int}}_{Q}$ increases as $\beta$ is raised. In a single Markov chain within HMC, once equilibrium is reached, the topological charge rarely evolves at large $\beta$, whereas the diffusion model exhibits significantly better ergodicity. DM also demonstrates stability across different lattice sizes, with $\tau^{\mathrm{int}}_{Q}\approx 130$ for all sizes at $\beta=7$. Additionally, we test a Metropolis-adjusted traditional Langevin sampler by replacing only the drift term in Algorithm 2 with $-\delta S[\phi]/\delta\phi(x,\tau)$, considering both single-chain and multi-chain settings. The results are unsatisfactory: for the single chain, $\tau^{\mathrm{int}}_{Q}$ remains significantly large even at small $\beta$, while for the multi-chain approach, the topological susceptibility is several times larger than the analytical result.
@@ Appendix C Analytical Results
[…]
where we used that the Haar measure with the defining representation Kanwar:2021wzm , $U=e^{-i\phi}$, is $d\phi/(2\pi)$. The integral is the modified Bessel function of the first kind, $I_{n=0}(\beta)$.
The partition function in fixed topology can also be derived from the known partition function in the $\theta-$vacuum Kovacs:1995nn ; Bonati:2019ylr ; Albandea:2021lvl . At sufficiently large volume ($V\rightarrow\infty$),
@@
[…]
In the second line we changed variables, which makes the integrated function decays rapidly beyond the interval $[-\pi,\pi]$. From the partition function, one can derive the probability density over $Q$,
[…]
@@
[…]
where $Q$ varies over all integers. These expressions are consistent with Refs. Kovacs:1995nn ; Bonati:2019ylr ; Bonati:2019olo ; Albandea:2021lvl .
On $L\times L$ lattices with periodic boundary conditions, the average value of the Wilson loops for U(1) gauge fields can be derived using integration over the set of plaquette variables after gauge fixing Kanwar:2021wzm ,
[…]
@@ Appendix D Likelihood Estimation
Samplers based on stochastic differential equations incorporate a noise term, making it impossible to directly estimate the likelihood of generated samples. In contrast, the probability flow of an ordinary differential equation (ODE) song2020score enables such likelihood estimation, as introduced in Ref. Wang:2023exq . The ODE sampler is sufficient for generating configurations at $\beta=1$. By comparing the log-likelihood of samples generated by the trained diffusion model with the physical action, we can assess whether the model has successfully captured the underlying physical distribution.

### v2

@@ Appendix A Langevin Sampler
4:  $\alpha_{i}\leftarrow\epsilon\cdot g\cdot t_{i}^{2}/t_{N_{T}}^{2}$ $\triangleright$ $\alpha_{i}$ is the step size.
@@
4:  $\alpha_{i}\leftarrow\epsilon\cdot g\cdot t_{i}^{2}/t_{N_{T}}^{2}$ $\triangleright$ $\alpha_{i}$ is the step size.
@@
9:   if $t_{i}<0.001$ then
10:     $P_{i}\leftarrow\min\left\{1,\frac{p(\tilde{\psi}_{\tau})q(\tilde{\phi}_{\tau-1}|\tilde{\psi}_{\tau})}{p(\tilde{\phi}_{\tau-1})q(\tilde{\psi}_{\tau}|{\tilde{\phi}_{\tau-1}})}\right\}$
@@ Appendix B Autocorrelation Time
Although in our approach multiple Markov chains are employed to generate independent samples, the dynamics in a single chain can be compared with other Monte Carlo methods. For traditional Monte Carlo methods, the characteristic length of autocorrelations for any observable $\mathcal{O}$ can be defined by the integrated autocorrelation time $\tau^{\mathrm{int}}_{\mathcal{O}}$ Madras and Sokal (1988).
@@
Fig. 7 compares the integrated autocorrelation time of the topological charge, $\tau^{\mathrm{int}}_{Q}$, for HMC and DM. For both methods, $\tau^{\mathrm{int}}_{Q}$ increases as $\beta$ is raised. In a single Markov chain within HMC, once equilibrium is reached, the topological charge rarely evolves at large $\beta$, whereas the diffusion model exhibits significantly better ergodicity. DM also demonstrates stability across different lattice sizes, with $\tau^{\mathrm{int}}_{Q}\approx 130$ for all sizes at $\beta=7$. Additionally, we test a Metropolis-adjusted traditional Langevin sampler by replacing only the drift term in Algorithm 2 with $\delta S[\phi]/\delta\phi(x,\tau)$, considering both single-chain and multi-chain settings. The results are unsatisfactory: for the single chain, $\tau^{\mathrm{int}}_{Q}$ remains significantly large even at small $\beta$, while for the multi-chain approach, the topological susceptibility is several times larger than the analytical result.
@@ Appendix C Analytical Results
$Z=\int dU_{\Box}\,e^{-S(U_{\Box})}=\int\prod_{\Box}dU\,e^{\frac{\beta}{2}[U+U^{-1}]}=\prod_{\Box}\int\frac{d\phi}{2\pi}\,e^{\beta\cos{\phi}}=[I_{0}(\beta)]^{V},$ (26)
where we used that the Haar measure with the defining representation Kanwar (2021), $U=e^{i\phi}$, is $d\phi/(2\pi)$. The integral is the modified Bessel function of the first kind, $I_{n=0}(\beta)$.
The partition function in fixed topology can also be derived from the known partition function in the $\theta$-vacuum Kovacs et al. (1995); Bonati and Rossi (2019b); Albandea et al. (2021). At sufficiently large volume ($V\rightarrow\infty$),
@@
$=\int_{-\infty}^{\infty}dk\,e^{-ikQ}\prod_{\Box}I_{k/(2\pi)}(\beta)=\int_{-\pi}^{\pi}d\theta\,e^{-i\theta Q}[I_{\theta/(2\pi)}(\beta)]^{V}.$ (27)
In the second line we changed variables, which makes the integrated function decays rapidly beyond the interval $\pi,\pi]$. From the partition function, one can derive the probability density over $Q$,
$p(Q)\equiv\frac{Z_{Q}}{Z}=\int_{-\pi}^{\pi}d\theta\,e^{-i\theta Q}\left[\frac{I_{\theta/(2\pi)}(\beta)}{I_{0}(\beta)}\right]^{V}.$ (28)
@@
$\chi_{Q}=\frac{1}{V}\int dQ\,p(Q)Q^{2},$ (29)
where $Q$ varies over all integers. These expressions are consistent with Refs. Kovacs et al. (1995); Bonati and Rossi (2019b, a); Albandea et al. (2021).
On $L\times L$ lattices with periodic boundary conditions, the average value of the Wilson loops for U(1) gauge fields can be derived using integration over the set of plaquette variables after gauge fixing Kanwar (2021),
$\langle W_{A}\rangle=\frac{1}{Z}\sum_{k}I_{k}(\beta)^{V-A}\left(\int dU_{\Box}\,U^{k+1}e^{\frac{\beta}{2}[U+U^{-1}]}\right)=\frac{\sum_{k}I_{k}(\beta)^{V-A}I_{k+1}(\beta)^{A}}{\sum_{k}I_{k}(\beta)^{V}},$ (30)
@@ Appendix D Likelihood Estimation
Samplers based on stochastic differential equations incorporate a noise term, making it impossible to directly estimate the likelihood of generated samples. In contrast, the probability flow of an ordinary differential equation (ODE) Song et al. (2020) enables such likelihood estimation, as introduced in Ref. Wang et al. (2024). The ODE sampler is sufficient for generating configurations at $\beta=1$. By comparing the log-likelihood of samples generated by the trained diffusion model with the physical action, we can assess whether the model has successfully captured the underlying physical distribution.
@@
Table 5: Comparison of the topological susceptibility and the $1\times 1$ Wilson loop for $L=16$ and $\beta=1,7,11$ using various algorithms.
coupling ($\beta$) Topological Susceptibility
p-HMC-Cold p-HMC-Hot HMC DM MALA-Hot MALA-Cold Exact
1 0.0406(19) 0.0398(18) 0.0416(16) 0.0422(17) 0.0421(20) 0.0404(18) 0.0406
7 0.0008(1) 0.0095(4) 0.00013(2) 0.0063(3) 0.0131(5) 0 0.0039
11 0 0.0148(6) 0 0.0058(3) 0.0165(13) 0 0.0024
coupling ($\beta$) $1\times 1$ Wilson Loop
p-HMC-Cold p-HMC-Hot HMC DM MALA-Hot MALA-Cold Exact
1 0.445(38) 0.4478(38) 0.447(37) 0.446(37) 0.444(36) 0.444(37) 0.446
7 0.927(7) 0.925(7) 0.926(7) 0.926(7) 0.924(6) 0.925(7) 0.926
11 0.954(4) 0.953(5) 0.954(3) 0.953(4) 0.950(5) 0.952(4) 0.953
Appendix E Different Sampler Results
Here we provide a more apples-to-apples comparison between the various algorithms. In Table 5 we show results for the topological susceptibility and the $1\times 1$ Wilson loop. We compare:
  • parallel HMC (p-HMC) with cold/hot starts (cold means the initial state is all zeroes; hot means the initial state is drawn from a standard Gaussian distribution);
  • single-chain hybrid Monte Carlo (HMC);
  • diffusion model (DM), including accept/reject and annealing. The reported DM results are the ‘ensemble-averaged’ outcomes: the mean over the models from the last eight training steps, each evaluated on 128 configurations. One could also consider Bayesian model averaging as an better alternative.;
  • Langevin dynamics including the accept/reject step (MALA), using cold/hot starts, all other setting is consistent with DM sampler in Appendix A except that the learned drift is replaced by the gradient of the target action.
For HMC and MALA a burn-in stage is always included before Metropolis is employed. Since at large $\beta$ every chain in HMC and MALA is frozen, the final result will depend more and more on the initial state. Hence for a hot start, the topological susceptibility at $\beta=11$ is larger than at $\beta=7$. The diffusion model is not sensitive to the initial state and topological sectors can be explored sufficiently well.